# On Temperature Scaling and Conformal Prediction of Deep Classifiers

### v3

@@ Abstract
In many classification applications, the prediction of a deep neural network (DNN) based classifier needs to be accompanied by some confidence indication. Two popular approaches for that aim are: 1) Calibration: modifies the classifier’s softmax values such that the maximal value better estimates the correctness probability; and 2) Conformal Prediction (CP): produces a prediction set of candidate labels that contains the true label with a user-specified probability, guaranteeing marginal coverage but not, e.g., per class coverage. In practice, both types of indications are desirable, yet, so far the interplay between them has not been investigated. Focusing on the ubiquitous Temperature Scaling (TS) calibration, we start this paper with an extensive empirical study of its effect on prominent CP methods. We show that while TS calibration improves the class-conditional coverage of adaptive CP methods, surprisingly, it negatively affects their prediction set sizes. Motivated by this behavior, we explore the effect of TS on CP beyond its calibration application and reveal an intriguing trend under which it allows to trade prediction set size and conditional coverage of adaptive CP methods. Then, we establish a mathematical theory that explains the entire non-monotonic trend. Finally, based on our experiments and theory, we offer simple guidelines for practitioners to effectively combine adaptive CP with calibration.
@@ 1 Introduction
Clearly, in critical applications, both calibration and CP are desirable, as they provide complementary types of information that can lead to a comprehensive decision. However, as far as we know, so far the interplay between them has not been investigated. Specifically, the works (Angelopoulos et al., 2021; Lu et al., 2022; Gibbs et al., 2023; Lu et al., 2023) apply initial TS calibration (rather than any other calibration method) before applying their CP methods. Yet, none of them investigates what is the effect of this procedure on the CP methods.
@@ 2.2 Conformal Prediction
Note that all these CP methods can be readily applied on $\bm{\pi}_{T^{*}}(\cdot)$ after TS calibration. This is done, in (Angelopoulos et al., 2021; Lu et al., 2022; Gibbs et al., 2023; Lu et al., 2023), where the authors stated that they applied TS calibration before examining the CP techniques. However, none of these works has examined how TS calibration impacts any CP method. All the more so, no existing work has experimented applying TS with a range of temperatures before employing CP methods.
@@ 3.1 Experimental setup
Metrics. We report metrics over the validation set samples that were not included in the calibration set or CP set. The metrics are as follows:
$\bullet$ Average set size (AvgSize) - The mean prediction set size of the CP algorithm. (See (16) for the definition.)
$\bullet$ Marginal coverage gap (MarCovGap) - The deviation of the marginal coverage from the desired $1-\alpha$. (See (17) for the definition.)
$\bullet$ Top-5% class-coverage gap (TopCovGap) - The deviation from the desired $1-\alpha$ coverage, averaged over the 5% of classes with the highest deviation. We use top-5% classes deviation due to the high variance in the maximal class deviation. (See (18) for the definition.)
The mathematical definitions of the metrics, along with additional details about the experimental setup, are presented in Appendix B.4. Note that for these metrics: the lower the better. Similar metrics have been used in (Ding et al., 2023; Angelopoulos et al., 2021). All the reported results, per metric, are the median-of-means along 100 trials where we randomly select the calibration/CP sets, similarly to (Angelopoulos et al., 2021).
@@ 4.2 The effect of TS on prediction set sizes of APS
Let $\hat{q}_{T}$ and $\hat{q}$ denote the thresholds of APS when applied with and without TS, respectively. For a new test sample $\mathbf{x}$ with logits vector $\mathbf{z}$ that is sorted in descending order, $\hat{\bm{\pi}}_{T}=\bm{\sigma}(\mathbf{z}/T)$ and $\hat{\bm{\pi}}=\bm{\sigma}(\mathbf{z})$ are the softmax outputs with and without TS, which are sorted as well. We denote $L_{T}=\min\{l:\sum\nolimits_{i=1}^{l}\hat{\pi}_{T,i}\geq\hat{q}_{T}\}$, $L=\min\{l:\sum\nolimits_{i=1}^{l}\hat{\pi}_{i}\geq\hat{q}\}$ as the prediction set sizes for this sample according to APS with and without TS, respectively.
@@
$&\text{If}\quad 0<T<1,\quad\text{then}:\\ &\forall M\in[L]:\quad\sum\nolimits_{i=1}^{M}\hat{\pi}_{i}-\hat{\pi}_{T,i}\leq\hat{q}-\hat{q}_{T}\implies L\geq L_{T}\\ &\text{If}\quad T>1,\quad\text{then}:\\ &\forall M\in[L_{T}]:\quad\sum\nolimits_{i=1}^{M}\hat{\pi}_{i}-\hat{\pi}_{T,i}\geq\hat{q}-\hat{q}_{T}\implies L\leq L_{T}\end{aligned}\right.$ (3)
@@
Let $\mathbf{z}^{q}$ and $\bm{\pi}^{q}$ denote the sorted logits vector and the softmax vector associated with the sample associated with APS threshold (without applying TS), that we dub “the quantile sample”. The CP theory builds on the quantile sample being larger than $(1-\alpha)\%$ of the scores of other samples with high probability.
Figure 2: $\Delta z$ for each sample sorted by score value (from low to high) with and without TS calibration, for CIFAR100-ResNet50.
As illustrated in Figure 2, there is a strong correlation between the score value and the difference $\Delta z:=z_{(1)}-z_{(2)}$. This implies that, for typical values of $\alpha$ (e.g., 0.1), the quantile sample exhibits a highly dominant first entry in its softmax vector, i.e., $\pi_{(1)}^{q}\gg\pi_{(2)}^{q}$ (recall that $\pi_{i}\propto\exp(z_{i})$). Similar behavior occurs with TS calibration, see Figure 2. Thus, if we denote $\mathbf{z}^{q}_{T}$ and $\bm{\pi}^{q}_{T}$ as the quantile sample when TS is used, we still have $\pi_{T,(1)}^{q}\gg\pi_{T,(2)}^{q}$. Consequently, $\bm{\pi}^{q}\approx\bm{\pi}^{q}_{T}$. Therefore, it is reasonable to make the technical assumption that both $\hat{q}$ and $\hat{q}_{T}$ correspond to the same sample in the CP set, denoted here by the sorted logits vector $\mathbf{z}^{q}$.
@@
With our assumption that $\hat{q}$ and $\hat{q}_{T}$ are associated with the same quantile sample $\mathbf{z}^{q}$, we have that $\hat{q}-\hat{q}_{T}$ in (3) can be written as $g(\mathbf{z}^{q};T,L^{q})$, where $L^{q}$ denotes the rank of the true label of $\mathbf{z}^{q}$. Empirically, we observed that $g(\mathbf{z}^{q},T,M)\approx g(\mathbf{z}^{q},T,L^{q})$ in our experiments (where $\Delta z^{q}=z^{q}_{(1)}-z^{q}_{(2)}\gg 1$). Furthermore, in Proposition A.6, we prove that $|g(\mathbf{z}^{q},T,M)-g(\mathbf{z}^{q},T,L^{q})|$ decays exponentially with $\Delta z^{q}$. See Appendix A.1 for more details. This justifies studying the following events:
@@
$\footnotesize b(T):=\left\{\begin{aligned} &\text{If}\quad 0<T<1:\\ &\max\left\{\frac{T}{T-1}\ln\left(\frac{T}{4}\right),\frac{T}{T+1}\ln\left(\frac{4(C-1)^{2}}{T}\right)\right\}\\ &\text{If}\quad T>1:\\ &\max\left\{\frac{T}{T-1}\ln(4T),\frac{T-1}{T+1}\ln(4T(C-1)^{2})\right\}\end{aligned}\right.$ (6)
@@
We now show that the bounds in Theorem 4.4 do not require unreasonable values of $\Delta z$ and $T$. In particular, the theorem explains the phenomenon for a “typical” $\mathbf{z}$ — e.g., the sample with the median score in Figure 2 — which in turn explains why we see the increase/decrease of the mean prediction set. Indeed, for this sample, according to Figure 2 we have $\Delta z\approx 8$. For $C=100$ (as in this CIFAR-100 experiment), the bounds in the theorem are complied by this sample for the temperature ranges $0.831\displaystyle 0<T<$ 0.831 and $4.81\displaystyle 1.25<T<$ 4.81. Since the calibrated temperature in this CIFAR100-ResNet50 experiment is $T^{*}=1.524$, which falls in this range, we rigorously proved increased prediction set for the median sample after TS calibration. Interestingly, the broad temperature range that is covered by our theory indicates that our bounds are sufficiently tight to establish additional fine-grained insights.
@@
We see that $\tilde{T}_{c}$ describes the temperature at which the trend of the prediction set size shifts from increasing to decreasing as $T$ increases. Our theory shows that $\tilde{T}_{c}$ shifts to lower temperatures as $C$ increases. A similar trend is observed in the empirical results shown in Figure 1. Specifically, for ImageNet-ViT ($C=1000$), the empirical maximum of the mean prediction set size occurs at $T_{c}=1.80$, while the bound’s minimum is at $\tilde{T}_{c}=1.78$. For CIFAR100-DenseNet121, the empirical maximum is at $T_{c}=2.32$ and the bound’s minimum at $\tilde{T}_{c}=2.12$. For CIFAR10-ResNet50, the empirical maximum is at $T_{c}=3.40$, and the bound’s minimum at $\tilde{T}_{c}=3.29$. Not only does our theoretical bound capture the non-monotonic trend, but it also provides $\tilde{T}_{c}$ close to the empirical $T_{c}$.
@@ 5 Guidelines for Practitioners
Based on our theoretically-backed findings, we propose a guideline, depicted in Fig. 4, for practitioners that wish to use adaptive CP methods (e.g., due to their better conditional coverage). Specifically, we suggest to use TS with two different temperature parameters on separate branches: $T^{*}$ that is optimized for TS calibration, and $\hat{T}$ that allows trading the prediction set sizes and conditional coverage properties of APS/RAPS to better fit the task’s requirements. Our experiments and theory show that $\hat{T}$ should be scanned up to a value $T_{c}$, which can be approximated by our theory.
@@
Note that one does not know in advance what values of the metrics are obtained per value of $\hat{T}$. However, since we propose to separate the calibration and the CP procedure, the calibration set can also be used to evaluate the CP algorithms without dangering exchangeability. Indeed, in Appendix C, we demonstrate how using a small amount of calibration data we can approximate the curves of AvgSize and TopCovGap vs. $T$ that appear in Figure 1 (which were generated using the entire validation set that is not accessible to the user in practice). According to the approximate trends, the user can choose $\hat{T}$ that best fit their requirements. Furthermore, note that the procedure required to produce approximated curves of metrics vs. $T$ is done offline during the calibration phase and its runtime is negligible compared to the offline training of DNNs.
In Appendix D, we further show the practical significance of our findings and guidelines. Specifically, for users that prioritize class-conditional coverage, we show that applying TS with temperature of the minimum in the approximated TopCovGap curve (created using only the calibration set as explained in Appendix C) followed by RAPS outperforms Mondrian CP (Vovk, 2012), which is tailored to classwise CP, in both TopCovGap and AvgSize metrics.
@@ Acknowledgment
The work is supported by ISF grant 1940/23 and by MOST grant 0007091.
@@
Since $z_{i}-z_{j}\geq 0$ and $\tilde{T}-\tilde{T}/T\geq 0$ we have that
$\exp\left[(z_{i}-z_{j})\left(\tilde{T}-\frac{\tilde{T}}{T}\right)\right]\geq 1,$
where the inequality is strict, unless $T=\tilde{T}$ or $z_{i}=z_{j}$. Next, observe that
$\displaystyle\exp\left[(z_{i}-z_{j})\left(\tilde{T}-\frac{\tilde{T}}{T}\right)\right]$ $\displaystyle=\exp\left[z_{i}\left(\tilde{T}-\frac{\tilde{T}}{T}\right)-z_{j}\left(\tilde{T}-\frac{\tilde{T}}{T}\right)\right]$
$\displaystyle=\frac{\exp\left(\frac{z_{i}}{\tilde{T}}+\frac{z_{j}}{T}\right)}{\exp\left(\frac{z_{i}}{T}+\frac{z_{j}}{\tilde{T}}\right)}$
$\displaystyle=\frac{\exp(z_{i}/\tilde{T})\cdot\exp(z_{j}/T)}{\exp(z_{i}/T)\cdot\exp(z_{j}/\tilde{T})}.$
Using the inequality we have $\frac{\exp(z_{i}/\tilde{T})\cdot\exp(z_{j}/T)}{\exp(z_{i}/T)\cdot\exp(z_{j}/\tilde{T})}\geq 1$, which concludes the proof of the lemma. ∎
@@
$&\text{If}\quad 0<T<1,\quad\text{then}:\quad\forall M\in[L]:\quad\sum\nolimits_{i=1}^{M}\hat{\pi}_{i}-\sum\nolimits_{i=1}^{M}\hat{\pi}_{T,i}\leq\hat{q}-\hat{q}_{T}\implies L\geq L_{T}\\ &\text{If}\quad T>1,\quad\text{then}:\quad\forall M\in[L_{T}]:\quad\sum\nolimits_{i=1}^{M}\hat{\pi}_{i}-\sum\nolimits_{i=1}^{M}\hat{\pi}_{T,i}\geq\hat{q}-\hat{q}_{T}\implies L\leq L_{T}\end{aligned}\right.$ (9)

### v4

@@ Abstract
In many classification applications, the prediction of a deep neural network (DNN) based classifier needs to be accompanied by some confidence indication. Two popular approaches for that aim are: 1) Calibration: modifies the classifier’s softmax values such that the maximal value better estimates the correctness probability; and 2) Conformal Prediction (CP): produces a prediction set of candidate labels that contains the true label with a user-specified probability, guaranteeing marginal coverage but not, e.g., per class coverage. In practice, both types of indications are desirable, yet, so far the interplay between them has not been investigated. Focusing on the ubiquitous Temperature Scaling (TS) calibration, we start this paper with an extensive empirical study of its effect on prominent CP methods. We show that while TS calibration improves the class-conditional coverage of adaptive CP methods, surprisingly, it negatively affects their prediction set sizes. Motivated by this behavior, we explore the effect of TS on CP beyond its calibration application and reveal an intriguing trend under which it allows to trade prediction set size and conditional coverage of adaptive CP methods. Then, we establish a mathematical theory that explains the entire non-monotonic trend. Finally, based on our experiments and theory, we offer guidelines for practitioners to effectively combine adaptive CP with calibration, aligned with user-defined goals.
@@ 1 Introduction
Clearly, in critical applications, both calibration and CP are desirable, as they provide complementary types of information that can lead to a comprehensive decision. However, so far the interplay between them remains largely unexplored. Specifically, the works (Angelopoulos et al., 2021; Lu et al., 2022; Gibbs et al., 2023; Lu et al., 2023) apply initial TS calibration (rather than any other calibration method) before applying their CP methods. Yet, none of them investigates what is the effect of this procedure on the CP methods.
@@ 2.2 Conformal Prediction
Note that all these CP methods can be readily applied on $\bm{\pi}_{T^{*}}(\cdot)$ after TS calibration. This is done, in (Angelopoulos et al., 2021; Lu et al., 2022; Gibbs et al., 2023; Lu et al., 2023), where the authors stated that they applied TS calibration before examining the CP techniques. However, none of these works has examined how TS calibration impacts any CP method. All the more so, no prior work has experimented applying TS with a range of temperatures before employing CP methods.
After completing our work, we became aware of a concurrent work (Xi et al., 2024), which also studies the effect of TS on CP methods. Yet, their results are only a small subset of our results. They do not consider conditional coverage and explore only a limited range of temperatures, which masks the non-monotonic effect on the prediction set size of APS and RAPS (which they do not compare to LAC). In contrast, our paper provides a complete picture of the effect of TS across a wide range of temperatures on both the prediction set size and the class-conditional coverage of APS, RAPS, and LAC. This complete picture shows practitioners that tuning the temperature for APS and RAPS introduces a trade-off effect, and we provide a practical way to control it. Lastly, note that several theoretical results on prediction sets and calibration have been reported in (Gupta et al., 2020). Yet, that work is limited to binary classification and does not provide an explanation for the fact that calibration affects CP methods differently and in a non-monotonic manner, as shown and analyzed in our paper.
@@ 3.1 Experimental setup
Metrics. We report metrics over the validation set, which we denote by $\{(\mathbf{x}_{i}^{(val)},y_{i}^{(val)})\}_{i=1}^{N_{val}}$, comprising of the samples that were not included in the calibration set or CP set. The metrics are as follows.
$\bullet$ Average set size (AvgSize) – The mean prediction set size of the CP algorithm:
$\text{AvgSize}=\frac{1}{N_{val}}\sum_{i=1}^{N_{val}}|\mathcal{C}(\mathbf{x}_{i}^{(val)})|.$
$\bullet$ Marginal coverage gap (MarCovGap) – The deviation of the marginal coverage from the desired $1-\alpha$:
[…]
$\bullet$ Top-5% class-coverage gap (TopCovGap) – The deviation from the desired $1-\alpha$ coverage, averaged over the 5% of classes with the highest deviation:
[…]
We also evaluate the average class-coverage gap (AvgCovGap) – a metric that measures the deviation from the target $1-\alpha$ coverage, averaged across all classes. The experimental results for this metric closely mirror those observed with the TopCovGap metric. Thus, we defer its formal definition and experimental evaluation to Appendix B.5.3.
Note that for these metrics: the lower the better. Similar metrics have been used in (Ding et al., 2023; Angelopoulos et al., 2021). All the reported results, per metric, are the median-of-means along 100 trials where we randomly select the calibration/CP sets, similarly to (Angelopoulos et al., 2021).
@@ 4.2 The effect of TS on prediction set sizes of APS
Let $\hat{q}_{T}$ and $\hat{q}$ denote the thresholds of APS when applied with and without TS, respectively. For a new test sample $\mathbf{x}$ with logits vector $\mathbf{z}$ that is sorted in descending order, $\hat{\bm{\pi}}_{T}=\bm{\sigma}(\mathbf{z}/T)$ and $\hat{\bm{\pi}}=\bm{\sigma}(\mathbf{z})$ are the softmax outputs with and without TS, which are sorted as well. We denote $L_{T}=\min\{\ell:\sum\nolimits_{i=1}^{\ell}\hat{\pi}_{T,i}\geq\hat{q}_{T}\}$, $L=\min\{\ell:\sum\nolimits_{i=1}^{\ell}\hat{\pi}_{i}\geq\hat{q}\}$ as the prediction set sizes for this sample according to APS with and without TS, respectively.
@@
Let $\mathbf{z}\in\mathbb{R}^{C}$ such that $z_{1}\geq z_{2}\geq\dots\geq z_{C}$, $\hat{\bm{\pi}}_{T}=\bm{\sigma}(\mathbf{z}/T)$ and $\hat{\bm{\pi}}=\bm{\sigma}(\mathbf{z})$. Let $\hat{q},\hat{q}_{T}\in(0,1]$ such that if $T>1$ then $\hat{q}\geq\hat{q}_{T}$ and if $1\displaystyle 0<T<$ 1 then $\hat{q}\leq\hat{q}_{T}$. Let $L=\min\{\ell:\sum\nolimits_{i=1}^{\ell}\hat{\pi}_{i}\geq\hat{q}\}$ , $L_{T}=\min\{\ell:\sum\nolimits_{i=1}^{\ell}\hat{\pi}_{T,i}\geq\hat{q}_{T}\}$. The following holds:
$&\text{If}\quad 0<T<1\text{ and }L_{T}>1\quad\text{then}:\\ \quad&\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{i}-\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{T,i}\leq\hat{q}-\hat{q}_{T}\implies L\geq L_{T}\\ &\text{If}\quad T>1\text{ and }L>1\quad\text{then}:\\ \quad&\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{i}-\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{T,i}\geq\hat{q}-\hat{q}_{T}\implies L\leq L_{T}\end{aligned}\right.$ (3)
Note that our definitions of $L$ and $L_{T}$ imply that their minimal value is 1. Thus, in the case of $1\displaystyle 0<T<$ 1 and $L_{T}=1$, we trivially have $L\geq L_{T}$. Similarly, in the case of $T>1$ and $L=1$, we trivially have $L\leq L_{T}$.
@@
Let $\mathbf{z}^{q}$ denote the logits vector of the sample associated with APS threshold (without applying TS), which we dub “the quantile sample”. The CP theory builds on the score of the quantile sample being larger than $(1-\alpha)\%$ of the scores of other samples with high probability.
Figure 2: $\Delta z$ for each sample sorted by APS score value (from low to high), with and without TS calibration applied before sorting the scores, for CIFAR100-ResNet50.
As illustrated in Figure 2, there is a strong correlation between the score value and the difference $\Delta z:=z_{(1)}-z_{(2)}$. This implies that, for typical values of $\alpha$ (e.g., 0.1), the quantile sample exhibits a highly dominant first entry in its logits vector $\mathbf{z}^{q}$.
Notably, even if the samples are sorted according to their scores after applying TS, this strong correlation still holds. Specifically, denoting by $\mathbf{z}^{q}_{T}$ the logits vector of the quantile sample when TS is applied, we observe in Figure 2 that it still exhibits a highly dominant first entry, with a $\Delta z$ value similar to that of $\mathbf{z}^{q}$. Recall that the softmax operation applied to the logits vectors is invariant to constant shifts and further amplifies the dominance of larger entries ($\pi_{i}\propto\exp(z_{i})$). Thus, having $z^{q}_{T,(1)}-z^{q}_{T,(2)}\approx z^{q}_{(1)}-z^{q}_{(2)}\gg 1$, practically implies that $\mathbf{z}^{q}_{T}\approx\mathbf{z}^{q}$ up to an additive constant. In Appendix B.6, we demonstrate this similarity across additional dataset–model pairs and temperature settings. Therefore, it is reasonable to make the technical assumption that both $\hat{q}$ and $\hat{q}_{T}$ correspond to the same sample in the CP set, denoted here by the sorted logits vector $\mathbf{z}^{q}$.
@@
With our assumption that $\hat{q}$ and $\hat{q}_{T}$ are associated with the same quantile sample $\mathbf{z}^{q}$, we have that $\hat{q}-\hat{q}_{T}$ in (3) can be written as $g(\mathbf{z}^{q};T,L^{q})$, where $L^{q}$ denotes the rank of the true label of $\mathbf{z}^{q}$.
In Proposition A.6 in Appendix A.1, we prove that $\forall M\in[C]$, $|g(\mathbf{z}^{q},T,L^{q})-g(\mathbf{z}^{q},T,M)|$ decays exponentially with $\Delta z$. Due to the large $\Delta z$ of the quantile sample illustrated in Figure 2, we approximate $g(\mathbf{z}^{q},T,L^{q})\approx g(\mathbf{z}^{q},T,M)$. Substituting $M=L_{T}-1$ or $M=L-1$, depending on the branch in (3), this justifies studying the following events:
@@
$\footnotesize b(T):=\left\{\begin{aligned} &\text{If}\quad 0<T<1:\\ &\max\left\{\frac{T}{T-1}\ln\left(\frac{T}{4}\right),\frac{T}{T+1}\ln\left(\frac{4(C-1)^{2}}{T}\right)\right\}\\ &\text{If}\quad T>1:\\ &\max\left\{\frac{T}{T-1}\ln(4T),\frac{{T}}{T+1}\ln(4T(C-1)^{2})\right\}\end{aligned}\right.$ (6)
@@
We now show that the bounds in Theorem 4.4 do not require unreasonable values of $\Delta z$ and $T$. In particular, the theorem explains the phenomenon for a “typical” $\mathbf{z}$ — e.g., the sample with the median score in Figure 2 — which in turn explains why we see the increase/decrease of the mean prediction set. Indeed, for this sample, according to Figure 2 we have $\Delta z\approx 8$. For $C=100$ (as in this CIFAR-100 experiment), the bounds in the theorem are complied by this sample for the temperature ranges $0.83\displaystyle 0<T<$ 0.83 and $1.25<T<{2.33}$. Since the calibrated temperature in this CIFAR100-ResNet50 experiment is $T^{*}=1.524$, which falls in this range, we rigorously proved increased prediction set for the median sample after TS calibration. Interestingly, the broad temperature range that is covered by our theory indicates that our bounds are sufficiently tight to establish additional fine-grained insights.
@@
We see that $\tilde{T}_{c}$ describes the temperature at which the trend of the prediction set size shifts from increasing to decreasing as $T$ increases. Our theory shows that $\tilde{T}_{c}$ shifts to lower temperatures as $C$ increases. A similar trend is observed in the empirical results shown in Figure 1.
@@ 5 Guidelines for Practitioners
Based on our theoretically-backed findings, we propose a guideline, depicted in Fig. 4, for practitioners that wish to use adaptive CP methods (e.g., due to their better conditional coverage). Specifically, we suggest to use TS with two different temperature parameters on separate branches: $T^{*}$ that is optimized for TS calibration, and $\hat{T}$ that allows trading the prediction set sizes and conditional coverage properties of APS/RAPS to better fit the task’s requirements. Our experiments and theory show that $\hat{T}$ should be scanned up to a value $T_{c}$.
@@
Note that one does not know in advance what values of the metrics are obtained per value of $\hat{T}$. However, since we propose to separate the calibration and the CP procedure, the calibration set can also be used to evaluate the CP algorithms without dangering exchangeability. Indeed, in Appendix C, we demonstrate how using a small amount of calibration data we can approximate the curves of AvgSize and TopCovGap vs. $T$ that appear in Figure 1 (which were generated using the entire validation set that is not accessible to the user in practice). According to the approximate trends, the user can choose $\hat{T}$ that best fit their requirements. Furthermore, note that the procedure required to produce approximated curves of metrics vs. $T$ is done offline during the calibration phase and its runtime is on the order of minutes.
In Appendix D, we further show the practical significance of our findings and guidelines. Specifically, for users that prioritize class-conditional coverage, we show that applying TS with temperature of the minimum in the approximated TopCovGap curve (created using only the calibration set as explained in Appendix C) followed by RAPS outperforms Mondrian CP (Vovk, 2012), which is tailored to classwise CP, in both TopCovGap and AvgSize metrics. Code for reproducing our experiments and applying our guidelines is available at https://github.com/lahavdabah/TS4CP.
@@ Acknowledgment
The work was supported by the Israel Science Foundation (No. 1940/23) and MOST (No. 0007091) grants.
Impact Statement
Our work focuses on investigating the interplay between approaches aiming at making modern classifiers reliable. Our finding can guide practitioners to use these approaches better. Improved reliability estimation can significantly influence and guide society toward a better future.
@@
Given that $z_{i}-z_{j}\geq 0$ and $T\geq\tilde{T}\geq 0$, it follows that:
$(z_{i}-z_{j})/\tilde{T}\geq(z_{i}-z_{j})/T.$
Since the exponential function is monotonically increasing, applying it to both sides of the inequality yields:
$\exp((z_{i}-z_{j})/\tilde{T})\geq\exp((z_{i}-z_{j})/T).$
Multiplying both sides by $\exp(z_{j}/T)\exp(z_{j}/\tilde{T})$, we obtain:
$\exp(z_{i}/\tilde{T})\cdot\exp(z_{j}/T)\geq\exp(z_{i}/T)\cdot\exp(z_{j}/\tilde{T}),$
which is the desired result. The strict monotonicity of the exponential function implies that equality holds only if $(z_{i}-z_{j})/\tilde{T}=(z_{i}-z_{j})/T$, i.e., if $T=\tilde{T}$ or $z_{i}=z_{j}$.
@@
$&\text{If}\quad 0<T<1\text{ and }L_{T}>1\quad\text{then}:\quad\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{i}-\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{T,i}\leq\hat{q}-\hat{q}_{T}\implies L\geq L_{T}\\ &\text{If}\quad T>1\text{ and }L>1\quad\text{then}:\quad\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{i}-\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{T,i}\geq\hat{q}-\hat{q}_{T}\implies L\leq L_{T}\end{aligned}\right.$ (9)